A man whose mass is 60 kg rides a Ferris wheel that has a radius of 20 m. The man has a scale available in the cart he is riding. If his speed during the ride is a constant 5 m/s, what would the scale read at top of the ride?
Answer #1

Ft = m * g-m*ac

Ft=m(g-(\frac{v^{2}}{r}))

h-60kg(9.8m/s2-( (5m/s)2 / (20n) ))

Ft - 513N

or

m- 52kg